OpenStudy (anonymous):

How many units apart are the points (5, 20) and (10, 8)?

OpenStudy (anonymous):

You'll need to use the distance formula for this. \[d=\sqrt{(x _1-x_2)^{2}+(y_1-y_2)^2}\] So 5-10=-5, square that and you get 25. 20-8=12, square that and you get 144. 25+144=169 Square root of 169 is 13. So d = 13.

OpenStudy (anonymous):

cool name ^^^, btw..:)

OpenStudy (anonymous):

Thanks. :D

OpenStudy (anonymous):

:)